OpenStudy (anonymous):

Integration using Table of integrals? Integrate ( e^x ) dx / (3-e^2x) I'm stumped...

OpenStudy (helder_edwin):

\[ \large \int\frac{e^x}{3-e^{2x}}\,dx \] ?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

on my table of integrals, I thought maybe using the int du/ u sqrt( a +bu) would work but idk how to set it up

OpenStudy (helder_edwin):

u already learned substitution, right?

OpenStudy (anonymous):

u sub yes... but im not sure for this one...

OpenStudy (anonymous):

we have to use the table of integrals...

OpenStudy (helder_edwin):

u use substitution in order to turn an integral that is not on the table into one that IS on the table.

OpenStudy (helder_edwin):

u can rewrite the integral into \[ \large \int\frac{e^x}{3-(e^x)^2}\,dx \] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so u would be e^x

OpenStudy (anonymous):

du is e^x as well?

OpenStudy (helder_edwin):

yes

OpenStudy (helder_edwin):

so if u=e^x then du=e^x dx therefore the integral becomes \[ \large \int\frac{1}{3-u^2}\,du \] agree?

OpenStudy (helder_edwin):

this one should be on your table. right?

OpenStudy (anonymous):

im looking...

OpenStudy (helder_edwin):

i have it. do u want me to post it?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

theres too many agh!

OpenStudy (helder_edwin):

no problem \[ \large \color{blue}{\int\frac{dx}{a^2-x^2}=\frac{1}{2a}\ln \left|\frac{a+x}{a-x}\right|+C} \] for \(a\neq0\)

OpenStudy (anonymous):

thank you so much!!!!

OpenStudy (helder_edwin):

compare with what we got \[ \large \int\frac{1}{3-u^2}\,du \] a^2=.....

OpenStudy (anonymous):

did you get this from a table? what number was it? i cant find this one in my table... but its definetly the right one i just cant find it

OpenStudy (helder_edwin):

we must have different tables. i have a russian book with me.

OpenStudy (helder_edwin):

what is the value of a^2 ?

OpenStudy (anonymous):

i see... ill keep looking..

OpenStudy (anonymous):

found it!

OpenStudy (anonymous):

I have another question regarding these integrals using the tables. could you help?

OpenStudy (anonymous):

ill post a new question so you get credit

OpenStudy (helder_edwin):

sure. but, shouldn't we first finish this one?

OpenStudy (anonymous):

i finished it i just needed to find the formula from the tables... a is sqrt 3 and then you simply plug the rest in right?

OpenStudy (helder_edwin):

and replace the value of u=e^x

OpenStudy (anonymous):

yes.

OpenStudy (helder_edwin):

ok.

OpenStudy (anonymous):

i will close this and open a new question..